# Probing Pre-trained Language Models on Code Changes: Insights from ReDef, a High-Confidence Just-in-Time Defect Prediction Dataset

### v1

@@ 3.2. PLM Evaluation Protocol
The first three variants serve as heuristic baselines, representing straightforward ways of exposing code modifications to PLMs without relying on prior proposals. These five encodings let us quantify the trade-off between context and change signal under realistic token budgets and identify the most effective representation for defect prediction.
In Stage 2, we applied controlled perturbations to test whether models behave consistently when change signals are deliberately altered. Each perturbation targeted the representation of change information while leaving the overall code form largely intact (Ribeiro et al., 2018; Jia and Liang, 2017).
Spurious change markers: In the After+Change markers format, the number of modified lines is first computed, then the same number of <CHG> tags are randomly inserted at arbitrary locations, injecting meaningless signals during both training and testing.
Swapped snapshots: Under the Before+After format, models are trained identically, but at test time the positions of function_before and function_after are swapped with 50% probability (e.g., “[Before]: function_after [After]: function_before”).
Reversed diff tags: Under the Diff with tags format, models are trained identically, but tested with the <ADD> and <DEL> tags inverted.
Swapped added/deleted blocks: In the AddedDeleted blocks format, the two sections are randomly swapped with 50% probability, so that “[ADDED LINES]” contains deletions and “[DELETED LINES]” contains additions during both training and testing.
These counterfactual variants are designed to reduce model performance: if a PLM truly understands the relationship between code before and after—that is, the semantics of the modification—then disrupting this information should lead to a drop in performance compared to the corresponding Stage 1 results. Conversely, if performance remains largely unchanged, it suggests that the model has not learned meaningful change semantics and is instead relying on superficial cues from the overall code context.
@@ 4.1. Research Questions
RQ2: Which input encodings yield the best PLM performance?
RQ3: Does PLM performance degrade under counterfactual perturbations of change representations?
@@ 4.2. Models and Baselines
To evaluate how PLMs capture the semantics of code changes, we experimented with the following pre-trained models. CodeT5+ and CodeBERT were used with their default maximum input limit of 512 tokens. UniXcoder uniquely supports longer contexts, so we evaluated it under both 512- and 1,024-token limit settings to test performance under constrained and full-capacity conditions.
Pre-trained Language Models.
CodeBERT. (Feng et al., 2020) We included CodeBERT, a 125M parameter RoBERTa-based encoder pre-trained for code understanding. Despite being one of the earlier PLMs, it has consistently shown strong and robust performance in defect prediction and classification tasks, making it an essential baseline for fair comparison (Xiao et al., 2023; Dahou et al., 2025).
CodeT5+. (Wang et al., 2023b) We used the 220M encoder-only variant, as detection tasks do not require generation capabilities. CodeT5+ has achieved state-of-the-art results across a wide range of software engineering tasks (Niu et al., 2023; Gao et al., 2024), making it a competitive and modern baseline.
@@
In addition to these PLMs, we implemented input encodings inspired by DeepJIT (diff with <ADD>/<DEL> tags) and JIT-Fine (addeddeleted block format). These were not used as standalone baselines, but rather as representation strategies integrated into our Stage 1 evaluation protocol.
@@ 4.3. Evaluation Metrics
We evaluate model performance using five standard classification metrics: Accuracy, Precision, Recall, F1 (F1-score), and PR-AUC. Among these, F1 is adopted as the primary metric, as it balances precision and recall and is widely used for imbalanced datasets such as JIT-SDP (Hoang et al., 2019; Kamei et al., 2012; Hoang et al., 2020).
@@
PR-AUC is the area under the Precision–Recall curve, which provides a threshold-independent view of the trade-off between precision and recall across all operating points.
@@ 4.4. Implementation Details
For each PLM–encoding configuration, we fine-tuned the model ten times with different random seeds (12345–12354) and report the mean and standard deviation to reduce stochastic variance. All models were trained under consistent hyperparameters: a learning rate of , batch size 8, and the AdamW optimizer with decoupled weight decay (Loshchilov and Hutter, 2017). Gradient accumulation was applied to stabilize updates, and a linear learning-rate scheduler with a 10% warm-up ratio was used to avoid abrupt changes in the early stages. Each run lasted for 6 epochs.
The best checkpoint was selected based on validation F1 (our primary metric) and subsequently used for test reporting. To address class imbalance, we applied the standard weighting scheme, assigning the positive class a weight of (King and Zeng, 2001; Lin et al., 2017).
All experiments were conducted on Tesla V100-PCIE-32GB GPUs in a shared academic cluster. Due to CPU bottlenecks and scheduling constraints, full GPU capacity was not always utilized. Training time was consistent across encoding strategies since only the input format varied; for reference, a single CodeBERT run at the 512-token limit required approximately 30 minutes on one GPU.
@@
For RQ3, each perturbation was paired with its corresponding original encoding and evaluated across the same 10 seeds. We tested the directional hypothesis that perturbations degrade performance using one-sided paired -tests (), complemented by Wilcoxon signed-rank tests (Wilcoxon, 1945). We also report paired effect size as Cohen’s and non-parametric percentile 95% bootstrap confidence intervals (CIs) for the mean difference using 10,000 resamples (Tibshirani and Efron, 1993). -values were adjusted for multiple comparisons using Holm’s method within each set of tests (e.g., across the five perturbations).
Following common conventions, we consider results statistically significant at . Effect sizes are interpreted using standard thresholds: partial (small), (medium), (large); Cohen’s (small), (medium), (large) (Cohen, 2013).
@@ 5. Experimental results
RQ2: Which input encodings yield the best PLM performance?
Table 3 summarizes the performance of all PLM–encoding combinations. Overall, Diff w/ tags and AddedDeleted consistently achieved the highest F1 across models, whereas After+Markers and Before+After performed worst. In contrast, these latter two encodings often ranked highest in terms of accuracy. This discrepancy is explained by class imbalance: ReDef contains over three times more clean (negative) than defective (positive) modifications, so encodings biased toward the majority class can appear more accurate while sacrificing F1. To compare models more rigorously beyond such raw patterns, we applied statistical tests as follows.
| Model | Encoding | F1 | PR-AUC | Accuracy | Precision | Recall |
@@
| AddedDeleted |
We conducted a two-way repeated-measures ANOVA with model (4 PLMs) and encoding (5 variants) as within-subject factors (Table 4). Results showed significant main effects of both model () and encoding (), both indicating large effect sizes. No interaction was found (), suggesting that the relative ranking of encodings was consistent across PLMs. Post-hoc pairwise tests with Holm correction revealed that Diff w/ tags and AddedDeleted significantly outperformed After+Markers and Before+After (all ), while performing comparably to each other. After-only showed intermediate performance, significantly better than some weaker encodings but worse than the top two in most cases. The pattern was consistent for CodeBERT, CodeT5+, and UniXcoder-512, whereas for UniXcoder-1,024 the differences diminished and were not statistically significant.
@@
| Model | 9.96 | 3 | 27 | 0.53 | 1.05 | |
| Encoding | 13.80 | 4 | 36 | 0.61 | 1.24 | |
| ModelEncoding | 0.82 | 12 | 108 | .63 | 0.08 | 0.30 |
@@
Across models, this advantage of change-focused encodings can be explained by two factors. First, as shown in Figure 3, only about 30–40% of functions fit within 512 tokens, so encodings that pass the entire function (After-only, After+Markers, Before+After) often truncate important edits. Second, supplying the entire function can also introduce excessive, mostly unchanged context that distracts the model from the actual modification. The weakness of Before+After illustrates both issues: two full functions must fit under the same budget, increasing truncation risk and noise. Supporting the second explanation, UniXcoder performed worse at 1,024 tokens than at 512, suggesting that longer sequences add distracting context rather than improving focus. Finally, in our setting which primarily targets code understanding, CodeBERT again emerged as the strongest PLM, in line with recent studies showing its robustness on classification and comprehension tasks (Xiao et al., 2023; Dahou et al., 2025).
@@
RQ3: Does PLM performance degrade under counterfactual perturbations of change representations? As shown in Table 5, across the four perturbations—Spurious change markers, Swapped snapshots, Reversed diff tags, and Swapped added/deleted blocks—performance remained effectively unchanged. Even when before/after alignment or polarity direction was disrupted, F1 differences stayed below 1%. This stability is counterintuitive: if PLMs were truly leveraging semantic signals from code changes, such distortions should have produced clear drops in performance metrics. Instead, the negligible differences suggest that the models are not semantically sensitive to modification content, but rather exploit surface-level regularities (e.g., token distributions, positional patterns) that persist under perturbation. In other words, the robustness observed here reflects insensitivity rather than genuine understanding.
@@
Statistical tests (paired t-tests, Wilcoxon signed-rank tests, bootstrap with Holm correction) confirmed this robustness. Only Reversed diff tags showed statistical significance ( after correction), but the absolute effect size was negligible (, Cohen’s ). In practice, such a tiny change does not indicate genuine semantic sensitivity. Taken together, these results reveal an important limitation: current PLMs achieve stable performance not because they robustly model code-change semantics, but because they fall back on shallow cues that remain unaffected by perturbations. This highlights a fundamental gap between snapshot-level code understanding and true modification reasoning, and points to the need for edit-aware pre-training objectives or architectures explicitly designed to capture change semantics.

### v2

@@ 3.2. Code Language Models Evaluation Protocol
The first three variants serve as heuristic baselines, representing straightforward ways of exposing code modifications to CLMs without relying on prior proposals. These five encodings let us quantify the trade-off between context and change signal under realistic token budgets and identify the most effective representation for defect prediction.
In Stage 2, we applied controlled perturbations to test whether models behave consistently when change signals are deliberately altered. Rather than simulating realistic noise, these serve as diagnostic counterfactuals to determine whether a model’s performance stems from a genuine understanding of code-change semantics (e.g., diff polarity and before/after temporal order) or from an over-reliance on superficial code features. To illustrate, consider a clean bug-fix that changes a risky null-pointer access use(p); to a safe guarded access if (p) use(p);. In the Swapped added/deleted blocks strategy, the representation is transformed as follows:
| Example: Semantic shift in a null-pointer check fix |
| Original (Clean): [DELETED LINES] use(p); [ADDED LINES] if (p) use(p); |
| A logical fix for a potential crash. |
| Perturbed (Clean): [DELETED LINES] if (p) use(p); [ADDED LINES] use(p); |
| A logical reintroduction of a bug (regression). |
If a model truly understands the semantics of a “fix,” its confidence in a “clean” label should significantly drop when presented with the perturbed version, which now logically represents a defect. A model whose predictions remain invariant to such reversals is likely relying on the presence of certain keywords (e.g., if, NULL) regardless of their roles in the change, making it unreliable for real-world deployment where distinguishing between a fix and a regression is critical.
We applied one specific perturbation to each of the four change-aware encodings introduced in Stage 1, excluding the After-only baseline as it lacks explicit change signals to disrupt. We intentionally varied the scope of these perturbations—applying them either across both training and testing or exclusively at test time—to explore two distinct diagnostic scenarios:
Spurious change markers: In the After+Change markers format, the number of modified lines is first computed, then the same number of <CHG> tags are randomly inserted at arbitrary locations. This injection of meaningless signals occurs during both training and testing.
Swapped snapshots: Under the Before+After format, models are trained on original data but perturbed only at test time by swapping the positions of function_before and function_after with a 50% probability (e.g., “[Before]: function_after [After]: function_before”).
Reversed diff tags: Under the Diff with tags format, models are trained on original data but perturbed only at test time with the <ADD> and <DEL> tags consistently inverted.
Swapped added/deleted blocks: In the AddedDeleted blocks format, the two sections are randomly swapped with a 50% probability during both training and testing, such that “[ADDED LINES]” contains deletions and “[DELETED LINES]” contains additions.
Each perturbation targeted the representation of change information while leaving the overall code form largely intact (Ribeiro et al., 2018; Jia and Liang, 2017). These counterfactual variants are designed to reduce model performance: if a CLM truly understands the relationship between code before and after—that is, the semantics of the modification—then disrupting this information should lead to a drop in performance compared to the corresponding Stage 1 results. Conversely, if performance remains largely unchanged, it suggests that the model has not learned meaningful change semantics and is instead relying on superficial cues from the overall code context.
@@ 4.1. Research Questions
RQ2: Which input encodings yield the best CLM performance?
RQ3: Does CLM performance degrade under counterfactual perturbations of change representations?
@@ 4.2. Models and Baselines
To evaluate how CLMs capture the semantics of code changes, we employed four representative models across distinct architectures. We included established encoder-based models—CodeBERT, CodeT5+, and UniXcoder—which have consistently demonstrated strong performance in JIT-SDP tasks (Jiang et al., 2025; Lin et al., 2023). Additionally, we incorporated Qwen2.5, a state-of-the-art decoder-based model. To the best of our knowledge, this study represents the first evaluation of Qwen2.5 within the JIT-SDP domain, despite its reported success in other code-related tasks.
Regarding input constraints, CodeBERT and CodeT5+ were evaluated using their default maximum limit of 512 tokens. In contrast, UniXcoder and Qwen2.5 were tested at both 512- and 1,024-token scales. For Qwen2.5 specifically, we extended the total token budget by 100 tokens (setting the limits to 612 and 1,124) to accommodate the necessary system prompt and role definitions without truncating the actual code content, thereby ensuring a fair comparison of the effective code context.
Subject Models.
CodeBERT (Feng et al., 2020) We included CodeBERT, a 125M parameter RoBERTa-based encoder pre-trained for code understanding. Despite being one of the earlier code language models, it has consistently shown strong and robust performance in defect prediction and classification tasks, making it an essential baseline for fair comparison (Xiao et al., 2023; Dahou et al., 2025).
CodeT5+ (Wang et al., 2023) We used the 220M encoder-only variant, as detection tasks do not require generation capabilities. CodeT5+ has achieved state-of-the-art results across a wide range of software engineering tasks (Niu et al., 2023; Gao et al., 2024), making it a competitive and modern baseline.
@@
Qwen2.5 (Hui et al., 2024) We included Qwen2.5-7B-Instruct, representing the current state-of-the-art in open-source decoder-based models. We selected the 7B variant as the standard representative of large-scale models to evaluate its reasoning capabilities over code modifications. To further distinguish between the model’s inherent zero-shot reasoning and its task-specific adaptation, we evaluated Qwen2.5 in two settings: an inference-only (zero-shot) baseline and a fine-tuned version.
In addition to these CLMs, we implemented input encodings inspired by DeepJIT (diff with <ADD>/<DEL> tags) and JIT-Fine (addeddeleted block format). These were not used as standalone baselines, but rather as representation strategies integrated into our Stage 1 evaluation protocol.
@@ 4.3. Evaluation Metrics
We evaluate model performance using five standard classification metrics: Accuracy, Precision, Recall, F1-score(F1), and PR-AUC. Among these, F1 is adopted as the primary metric, as it balances precision and recall and is widely used for imbalanced datasets such as JIT-SDP (Hoang et al., 2019; Kamei et al., 2012; Hoang et al., 2020).
@@
PR-AUC is the area under the Precision–Recall curve, which provides a threshold-independent view of the trade-off between precision and recall across all operating points. Note that for the decoder-based model (Qwen2.5), we primarily report the four discrete classification metrics, as calculating threshold-independent measures like PR-AUC is less straightforward due to its generative inference nature compared to encoder-based models with dedicated classification heads.
@@ 4.4. Implementation Details
To ensure the reproducibility and statistical significance of our results, we fine-tuned each CLM–encoding configuration ten times using different random seeds (12345–12354) and report the mean and standard deviation. All experiments were conducted using NVIDIA Tesla V100 (32GB) and RTX 3080 Ti (12GB) GPUs.
Encoder-based Training. CodeBERT, CodeT5+, and UniXcoder were fine-tuned for 6 epochs with a learning rate of , a batch size of 8, and the AdamW optimizer with decoupled weight decay (Loshchilov and Hutter, 2017). We employed a linear learning-rate scheduler with a 10% warm-up ratio and utilized gradient accumulation to stabilize training. To address class imbalance, we applied a cost-sensitive weighting scheme, assigning the positive class a weight of (King and Zeng, 2001). The best-performing model across the 6 epochs was selected based on validation F1 for final testing.
Decoder-based Training. For Qwen2.5-7B-Instruct, we utilized the Unsloth (Daniel Han and team, 2023) framework to perform memory-efficient Supervised Fine-Tuning (SFT) with 4-bit quantization. Given the large parameter scale, we applied Parameter-Efficient Fine-Tuning (PEFT) via LoRA () to the query, key, value, and projection layers. The model was trained for a single epoch with a learning rate of , an effective batch size of 16, and the 8-bit AdamW optimizer. To ensure optimal convergence within a single epoch, we performed intra-epoch validation at 10 equally spaced intervals and selected the checkpoint with the highest validation F1. Unlike the encoder models, class imbalance was addressed via oversampling the defective class to match the size of the clean class (Buda et al., 2018). The same prompt templates were utilized for both fine-tuning and zero-shot evaluations to ensure consistency, all of which are available in our replication package.
@@
For RQ3, each perturbation was paired with its corresponding original encoding and evaluated across the same 10 seeds. We tested the directional hypothesis that perturbations degrade performance using one-sided paired -tests (), complemented by Wilcoxon signed-rank tests (Wilcoxon, 1945). We also report paired effect size as Cohen’s and non-parametric percentile 95% bootstrap confidence intervals (CIs) for the mean difference using 10,000 resamples (Tibshirani and Efron, 1993). -values were adjusted for multiple comparisons using Holm’s method within each set of tests (e.g., across the four perturbations). Following common conventions, we consider results statistically significant at . Effect sizes are interpreted using standard thresholds: partial (small), (medium), (large); Cohen’s (small), (medium), (large) (Cohen, 2013).
@@ 5. Experimental results
RQ2: Which input encodings yield the best CLM performance?
Table 4 summarizes the performance of all CLM–encoding combinations. Overall, Diff w/ tags and AddedDeleted consistently achieved the highest F1 across most models, while After+Markers and Before+After often performed worst, particularly in encoder-based models. Notably, we observed a recurring discrepancy where encodings with lower F1 often rank higher in accuracy. This discrepancy reflects the class imbalance in ReDef (3:1 clean-to-defective ratio), where accuracy is often inflated by a bias toward the majority clean class despite poor defect detection performance.
Beyond these task-specific adaptation results, Qwen2.5’s zero-shot performance exhibits a universal decline across all metrics. The After-only zero-shot configuration fails almost entirely (F1 0.1), underscoring that in zero-shot settings, providing explicit change information is crucial for the model to establish any predictive utility. To compare the fine-tuned models more rigorously beyond such raw patterns, we applied statistical tests as follows.
| Model | Encoding | F1 | PR-AUCa | Accuracy | Precision | Recall |
@@
| AddedDeleted | ||||||
| Qwen2.5-512 | After-only | — | ||||
| After+Markers | — | |||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| Qwen2.5-1,024 | After-only | — | ||||
| After+Markers | — | |||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| Qwen2.5-512 | After-only | — | ||||
| (zero-shot) | After+Markers | — | ||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| Qwen2.5-1,024 | After-only | — | ||||
| (zero-shot) | After+Markers | — | ||||
| Before+After | — | |||||
| Diff w/ tags | — | |||||
| AddedDeleted | — | |||||
| aPR-AUC is omitted for Qwen2.5 due to its generative output nature. | ||||||
We conducted a two-way repeated-measures ANOVA with model (6 fine-tuned CLMs) and encoding (5 variants) as within-subject factors (Table 5). Results showed significant main effects of both model () and encoding (), both indicating large effect sizes. Post-hoc pairwise tests with Holm correction on the marginal means (collapsed over models) confirmed that Diff w/ tags and AddedDeleted significantly outperformed all other encodings (all ), while performing comparably to each other.
@@
| Model | 9.32 | 5 | 45 | 0.51 | 1.02 | |
| Encoding | 15.11 | 4 | 36 | 0.63 | 1.30 | |
| ModelEncoding | 1.91 | 20 | 180 | .014 | 0.17 | 0.46 |
Furthermore, a significant interaction effect was found (), suggesting that the impact of encoding strategies varies across different model architectures. Simple effect analysis revealed that for CodeBERT and CodeT5+, change-focused encodings (Diff w/ tags, AddedDeleted) consistently demonstrated a statistically significant performance advantage (). In contrast, for UniXcoder and Qwen2.5 models, the pairwise differences did not reach statistical significance. Nevertheless, a consistent numerical advantage of change-focused encodings was observed across all models, with Diff w/ tags or AddedDeleted consistently achieving the highest F1 (Table 4).
@@
This consistent advantage can be explained by two primary factors: truncation risk and distracting context. First, as shown in Figure 3, only about 30–50% of functions fit within the 512-token limit. Consequently, whole-function encodings (After-only, After+Markers, Before+After) often truncate important edits, whereas compact diff-style formats preserve the critical change signal within the token budget. Second, supplying the entire function introduces excessive, mostly unchanged context that distracts the model from the actual modification. The weakness of Before+After illustrates both issues: two full functions must fit under the same budget, simultaneously increasing truncation risk and noise. Supporting this distracting context hypothesis, both UniXcoder and Qwen2.5 generally performed better under a 512-token limit than at 1,024 tokens. Therefore, the observed superiority of diff-style encodings is primarily a practical one; it reflects their efficiency in preserving critical signals within finite context windows rather than an inherent semantic advantage over whole-function views.
Interestingly, we observed an inverse correlation between model scale and performance. Despite having the largest parameter scale, Qwen2.5 exhibited the lowest overall performance. This trend extends to encoder-based models, where the model with the fewest parameters—CodeBERT—achieved the highest F1. Given that ReDef consists of difficult samples—modifications that professional developers initially committed and subsequently reverted—the signals for defects are likely extremely subtle. In such high-difficulty scenarios, larger models may struggle to discern meaningful features, while smaller models appear more adept at capturing simpler but consistent patterns.
While the absolute performance remains modest (F1 scores below 0.4), such results are common in classification tasks involving severe class imbalance and subtle signals, as characterized by the ReDef dataset (Guo et al., 2023; Ni et al., 2023). Nevertheless, as all models perform substantially above random baselines, they are evidently capturing discriminative patterns. This observation sets the stage for RQ3, where we investigate whether these learned patterns originate from a genuine understanding of code-change semantics or a reliance on shallow, invariant surface cues.
Finally, in our setting which primarily targets code understanding, CodeBERT again emerged as the strongest CLM, in line with recent studies showing its robustness on classification and comprehension tasks (Xiao et al., 2023; Dahou et al., 2025).
RQ3: Does CLM performance degrade under counterfactual perturbations of change representations?
@@
To address RQ3, we conducted counterfactual perturbations using CodeBERT, the best-performing encoder-based model, and Qwen2.5-512, the strongest decoder-based variant. As shown in Table 6, across the four perturbations—Spurious change markers, Swapped snapshots, Reversed diff tags, and Swapped blocks—performance for both CodeBERT and Qwen2.5-512 remained effectively stable. Even when the logical direction or structural alignment of the modification was disrupted, F1 differences stayed within a negligible range (). This stability is counterintuitive: if these models were truly leveraging semantic signals from code changes, such distortions should have produced clear drops in performance metrics. Instead, the negligible differences suggest that the models are not semantically sensitive to modification content, but rather exploit surface-level regularities (e.g., token distributions, positional patterns) that persist under perturbation. In other words, the robustness observed here reflects insensitivity rather than genuine understanding.
The lack of response in the Swapped snapshots and Swapped blocks conditions ( after Holm correction for both models) provides compelling evidence of semantic blindness. Logically, reversing the states or blocks represents a functional opposite, yet both models were impervious to these changes. One might attribute this stability to the modest baseline performance; however, even a model with limited predictive power should exhibit a clear drop when its underlying causal logic is fundamentally reversed. Despite this consistent trend of insensitivity, we observed that Qwen2.5-512 exhibited slightly more sensitivity in certain conditions. Specifically, it showed a statistically significant decline under Reversed diff tags () and a marginal drop under Spurious change markers (). In contrast, CodeBERT remained almost entirely unaffected by these same perturbations, suggesting that encoder-based architectures may be even more reliant on shallow, invariant cues than their decoder-based counterparts. However, even for the larger decoder-based model, the absolute magnitude of these drops remained remarkably small, further confirming that its predictive utility still relies heavily on surface-level regularities that persist under perturbation. These results highlight a critical semantic gap: current CLMs achieve stability not through robust modeling of code-change logic, but by falling back on shallow features. This underscores the urgent need for architectures or training objectives specifically designed to capture the relational dynamics of code transformations.
@@
| Qwen2.5-512 | After+Markers | — | ||||
| Spurious CHG markers | — | |||||
| Qwen2.5-512 | Before+After | — | ||||
| Swapped snapshots | — | |||||
| +0.30% | ||||||
| Qwen2.5-512 | Diff w/ tags | — | ||||
| Reversed diff tags | — | |||||
| Qwen2.5-512 | AddedDeleted | — | ||||
| Swapped blocks | — | |||||
| +0.17% |